OpenStudy (idealist10):

Solve (x^3)y'=2(y^2+x^2*y-x^4) explicitly.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

Do I multiply the right side?

OpenStudy (anonymous):

\[\begin{align*} x^3y'&=2y^2+2x^2y-2x^4\\\\ y'&=\frac{2y^2}{x^3}+\frac{2y}{x}-2x \end{align*}\] I'd try the homogeneous substitution next, \(y=vx\), then \(y'=xv'+v\). \[\begin{align*} xv'+v&=\frac{2v^2}{x}+2v-2x\\\\ v'&=\frac{2v^2}{x^2}+\frac{v}{x}-2 \end{align*}\] Another homogeneous sub should do the trick.

OpenStudy (idealist10):

After you solved for v', what would you do?

OpenStudy (anonymous):

Substitute again, \(v=tx\) so \(v'=xt'+t\). \[xt'+t=2t^2+t-2\] which will be a separable equation.

OpenStudy (idealist10):

I'll come back later due to something. Excuse me.

OpenStudy (idealist10):

Thank you!